# SAM-LAD: Segment Anything Model Meets Zero-Shot Logic Anomaly Detection

### v2

@@ Abstract
Visual anomaly detection is vital in real-world applications, such as industrial defect detection and medical diagnosis. However, most existing methods focus on local structural anomalies and fail to detect higher-level functional anomalies under logical conditions. Although recent studies have explored logical anomaly detection, they can only address simple anomalies like missing or addition and show poor generalizability due to being heavily data-driven. To fill this gap, we propose SAM-LAD, a zero-shot, plug-and-play framework for logical anomaly detection in any scene. First, we obtain a query image’s feature map using a pre-trained backbone. Simultaneously, we retrieve the reference images and their corresponding feature maps via the nearest neighbor search of the query image. Then, we introduce the Segment Anything Model (SAM) to obtain object masks of the query and reference images. Each object mask is multiplied with the entire image’s feature map to obtain object feature maps. Next, an Object Matching Model (OMM) is proposed to match objects in the query and reference images. To facilitate object matching, we further propose a Dynamic Channel Graph Attention (DCGA) module, treating each object as a keypoint and converting its feature maps into feature vectors. Finally, based on the object matching relations, an Anomaly Measurement Model (AMM) is proposed to detect objects with logical anomalies. Structural anomalies in the objects can also be detected. We validate our proposed SAM-LAD using various benchmarks, including industrial datasets (MVTec Loco AD, MVTec AD), and the logical dataset (DigitAnatomy). Extensive experimental results demonstrate that SAM-LAD outperforms existing SoTA methods, particularly in detecting logical anomalies.
Index Terms:
Anomaly detection, Anomaly localization, Zero-shot, Segment Anything Model, Keypoint Matching.
I Introduction
@@
In recent years, image anomaly detection techniques have been widely applied in industrial quality detection[1][2][3][4], anomaly segmentation[5], and medical diagnosis scenarios[6], aiming to detect abnormal data that are different from normal data within a sample image[7]. Since abnormal prior information is scarce, this task is commonly conducted within the framework of an unsupervised learning paradigm. Consequently, there has been increasing interest among scholars in researching unsupervised anomaly detection, and near-perfect results have been achieved, as evidenced by methods such as Pull&Push[8], PMB-AE[9], and Patchcore[10]. However, due to the limits of those scenarios, most anomaly detection methods currently focus on structural anomaly detection only can deal with even one object in a single scene at at a time, as illustrated in Fig. 1(a).
@@
In many usual scenarios such as autonomous driving and surveillance systems, understanding the semantic context of the entire scene and detecting anomalies is essential. Therefore, Bergmann et al.[11] have proposed logical anomalies, which represent more complex abnormalities in the logical relationships between objects within the entire scene. Specifically, within a scene, anomalies that appear normal at the local level but violate geometric constraints or logical principles when considering global semantics are referred to as global logical anomalies. For example, in Fig. 1(b), anomaly-free images of the category breakfast box always contain exactly two tangerines and one nectarine that are always located on the left-hand side of the box. Furthermore, the ratio and relative position of the cereals and the mix of banana chips and almonds on the right-hand side are fixed. An anomaly-free screw bag contains exactly two washers, two nuts, one longscrew, and one short screw. Nevertheless, their logical anomalies could involve a missing object, an unexpected object, or incorrect ordering and combination.
@@
For the more challenging global logical anomalies, most existing methods tailored for structural anomalies demonstrate catastrophic results, for instance, as illustrated in Fig. 1(c) with Patchcore. This is because structural or textural anomalies belong to lower-level anomaly types, which do not necessitate an understanding of the overall semantics of objects in the scene and only require local knowledge for anomaly detection. However, for higher-level global logical anomalies, relying solely on local perception to ascertain the normalcy of overall semantics is insufficient. Consequently, the performance of existing methods is significantly constrained.
@@
Our previous work[12] proposed a reconstruction-based method for detecting multiple object anomalies to address logical anomalies. In [12], objects had fixed spatial relationships. We used semi-supervised learning on positive samples to regress object positions in the test samples and reconstruct their normal features. However, in more complex scenarios like [11], where objects lack fixed spatial relationships, exploring new methods is essential to address logical anomalies.
@@
To address the above challenges: 1) multiple key objects in complex scenes. 2) variety of positive sample features. 3) poor generalization because of dependence on data-driven. We propose a novel framework called SAM-LAD, which can be plug-and-play in any scenario without training and even outperforms existing data-driven logical anomaly detection methods. Specifically, we ingeniously leverage the robust object segmentation capability of the Segment Anything Model (SAM)[16] to obtain the positional information of all key objects in the scene by setting segmentation thresholds. Then, we utilize a pre-trained DINOv2[17] backbone network as a feature extractor to extract features and employ nearest neighbor search to find the $k$ most similar normal samples to the query image. Following this, we implement the FeatUp[18] operation to upsample the feature maps and recover their lost spatial information. Combining the upsampled feature maps with the positional information obtained from SAM yields separate feature maps for each object. Subsequently, we propose a Dynamic Channel Graph Attention (DCGA) mechanism to effectively compress each object’s feature map into a single feature vector. Consequently, leveraging the proposed Object Matching Model (OMM), we match the feature vectors of the reference images with those of each feature vector in the query image. Finally, based on the matching results, we propose an Anomaly Measurement Model (AMM), which estimates the feature distribution of individual objects in the query image and the corresponding $k$ matched objects from the $k$ normal samples, thereby calculating the final anomaly score map. Noteworthy, thanks to the introduction of AMM, our framework not only excels in detecting global logical anomalies but also fulfills the requirements for detecting structural anomalies. Our framework is motivated by the following: when humans discern anomalies among multiple objects within intricate scenes, the most simplest and efficacious approach involves juxtaposing several normal samples with each object in the test sample individually. Through this meticulous comparison, inference regarding anomalous regions can be inferred without necessitating the establishment of a costly global semantic contextual comprehension within the model. The essence lies in identifying and aligning pivotal objects, thus accomplishing the entirety of the task. We evaluate the proposed framework on multiple commonly used benchmarks, and the experimental results demonstrate that our SAM-LAD achieves state-of-the-art performance. The main contributions of this paper can be summarized as follows:
@@
    We novelty introduce SAM into the field of logical anomaly detection, leveraging its strong segmentation prowess to obtain the positional information of each crucial object within complex scenes without training. With this positional information as the foundation, subsequent matching between objects can be facilitated.
@@
    We propose the DCGA mechanism, which, drawing from the principles of graph neural networks, can effectively compress the upsampled feature maps of individual objects while considering the logical dependencies among them, thereby preserving their most crucial features in the form of feature vectors. DCGA ensures the accuracy and efficiency of subsequent object matching.
@@
    We innovatively propose the OMM, which leverages the principles of optimal transport and the feature vectors obtained from DCGA to generate a matching score matrix between each pair of objects. Subsequently, utilizing the Sinkhorn algorithm to solve the matching score matrix yields the matching relationships between objects. Based on these relationships, anomalies within objects can be accurately identified.
  4. 4.
    We propose the AMM, which computes the feature distribution differences between a single object in the query image and the matched objects from positive samples. This model enables not only the detection of logical anomalies but also the identification of structural anomalies present within them.
II Related Work
II-A Semantic Segmentation
@@
II-B Keypoint Matching
Keypoint matching is essential in computer vision for object recognition, 3D reconstruction, and image stitching. The long-standing interest in this problem has led to various approaches. Early methods like SIFT[22] and ORB[23] laid the groundwork with handcrafted keypoint descriptors. These algorithms created robust descriptors invariant to scale, rotation, and partial occlusion, making them effective for many tasks despite sensitivity to image illumination changes. Recently, deep learning techniques revolutionized keypoint matching by learning discriminative features from large datasets. CNNs, like the SuperPoint model[24], have been widely adopted for keypoint detection and matching, learning detection and descriptor generation end-to-end. Deep learning approaches offer improved robustness and generalization over handcrafted methods. Recent advances use attention mechanisms to improve keypoint matching in challenging environments. Networks like SuperGlue[25] use graph neural networks with self- and cross-attention to establish robust keypoint correspondences. Transformers in keypoint matching improve context understanding, handling complex matching scenarios more proficiently.
@@ II-B Keypoint Matching
II-C Unsupervised Anomaly Detection
@@
Current unsupervised anomaly detection approaches are typically classified into two categories: methods targeting structural anomalies and those targeting logical anomalies. Most current anomaly detection methods target structural anomalies, using robust feature extraction networks to obtain high-level semantic features of a test image. The distance between features of test images and anomaly-free images is calculated, identifying areas with large distances as anomalies. Existing methods typically use deep CNN models like ResNet[26] and EfficientNet[27] for feature extraction. For example, Roth et al.[10] proposed Patchcore to detect anomalies on objects with a concise background. They selected ResNet-50[26] pretrained on ImageNet[28] as the feature extractor. However, these methods struggle with logical anomalies since the local structures within the image appear normal. Therefore, more researchers are focusing on logical anomaly detection. For instance, Zhang et al. proposed DSKD[15], and Batzner introduced EfficientAD[29]. Both methods use the Student-Teacher network, detecting anomalies by comparing teacher and student differences. Using a local-global branching approach, Yao et al. presented GLCF[13], while Bergmann et al. proposed GCAD[11] to understand the logical semantic constraints of the entire image scene. Guo et al. presented THFR[14], based on template-guided reconstruction. These methods have shown promise in detecting certain logical anomalies, such as missing components and unexpected excess. However, their performance drops significantly with more complex logical anomalies, such as misordering, mismatches, and haphazard object arrangements.
@@
The proposed SAM-LAD builds a zero-shot method without strenuous efforts to comprehend the logical semantics of the entire scene. By introducing SAM and proposing OMM and AMM, a global logical anomaly detection system was developed, effectively addressing the limitations of existing methods mentioned above.
III Methodology
@@
Figure 2: Overview of the proposed framework, which consists of 4 stages. The first stage is an offline operation, which is building an anomaly-free template features bank. The second stage is to extract a feature map from a query image. The next stage is to utilize SAM to further obtain object feature maps. The last stage involves matching the objects in the query image with those in the reference images one by one. Using the obtained matching relationships, we calculate anomalies within each object and obtain the final anomaly score maps.
III-A Architecture Overview
@@
The data flow of the proposed SAM-LAD is depicted in Fig. 2, which consists of four stages. 1) On-boarding stage 1 is an offline operation. For all normal images $\mathbb{R}^{3\times H\times W}$, a pre-trained backbone as a feature extractor is first deployed to extract the features from all normal images. Subsequently, the extracted features are compiled and stored within a template features bank. 2) Stage 2 involves extracting features from a query image $I_{q}\in\mathbb{R}^{3\times H\times W}$ using the same feature extractor as stage 1. 3) Stage 3, for the feature maps $f_{q}$ extracted from a query image $I_{q}$, we retrieve its $k$-nearest normal feature maps $(f_{r}^{i})_{i\in{[1,k]}}$ in the template bank and their corresponding images $(I_{r}^{i})_{i\in{[1,k]}}$. We designate the $k$ normal images procured as reference images and in parallel with the query image, input $(I_{q},I_{r}^{i})_{i\in{[1,k]}}$ into the SAM to obtain the individual object mask. Concurrently, the feature maps $(f_{q},f_{r}^{i})_{i\in{[1,k]}}$ are subjected to a FeatUp operation, upsampling to 8$\times$. By combining the individual object masks with the upsampled feature maps, we obtain the object feature maps for both the query and reference images, respectively. 4) In Stage 4, the DGCA module is employed to encode and compress the object feature maps into object descriptor vectors. Subsequently, utilizing the proposed Object Matching Module, the two sets of object descriptor vectors are matched. Based on the matching results, the Anomaly Measurement Module computed the ultimate anomaly score map.
III-B Feature Extraction and Template Features Bank
@@
The first and second stage of the proposed SAM-LAD is the extraction of feature maps. The same feature maps are later used for FeatUp operation. There are multiple options for extracting features. Recently, Vision Transformer (ViT) has exhibited remarkable performance in anomaly detection-related tasks[30] [31] due to its self-attention mechanism, enabling the model to attend to global information of the entire image when processing image patches. Therefore, we have adopted the pre-trained ViT-type DINOv2-S backbone[17], as the feature extractor for the proposed SAM-LAD. For a given image $I_{x}^{3\times H\times W}$, we denote the extracted feature maps $f_{x}^{C\times H^{{}^{\prime}}\times W^{{}^{\prime}}}$:
@@
At initialization, we execute offline operations to construct a template features bank =$\mathcal{B}=\left\{B_{1},\cdots,B_{i},\cdots,B_{Z},\right\}$ using $F$ to extract $Z$ normal images in the all normal set $\mathbb{R}^{3\times H\times W}$, where $B_{i}$ represents the template feature map of the $i$-th normal sample. At inference, only the feature maps of the query image are extracted and we could reduce the template features bank using coreset subsampling method[32][33] to reduce inference time and memory usage.
III-C Image-level Nearest Neighbor Search

### v4

@@ Abstract
Visual anomaly detection is vital in real-world applications, such as industrial defect detection and medical diagnosis. However, most existing methods focus on local structural anomalies and fail to detect higher-level functional anomalies under logical conditions. Although recent studies have explored logical anomaly detection, they can only address simple anomalies like missing or addition and show poor generalizability due to being heavily data-driven. To fill this gap, we propose SAM-LAD, a zero-shot, plug-and-play framework for anomaly detection in any scene. First, we obtain a query image’s feature map using a pre-trained backbone. Simultaneously, we retrieve the reference images and their corresponding feature maps via the nearest neighbor search. Then, we introduce the Segment Anything Model (SAM) to obtain object masks of the query and reference images. Each object mask is multiplied by the entire image’s feature map to obtain object feature maps. Next, an Object Matching Model (OMM) is proposed to match objects in the query and reference images. To facilitate object matching, we propose a Dynamic Channel Graph Attention (DCGA) module, treating each object as a keypoint and converting its feature maps into feature vectors. Finally, based on the object matching relations, an Anomaly Measurement Model (AMM) is proposed to detect objects with logical anomalies. Structural anomalies in the objects can also be detected. We validate our proposed SAM-LAD using various benchmarks, including industrial datasets (MVTec Loco AD, MVTec AD), and the logical dataset (DigitAnatomy). Extensive experimental results demonstrate that SAM-LAD outperforms existing SoTA methods, particularly in detecting logical anomalies.
keywords:
Anomaly detection, Anomaly localization, Zero-shot, Segment Anything Model, Keypoint Matching
journal: Knowledge-Based Systems
[…]
[label1]organization= Shanghai Institute of Intelligent Science and Technology, College of Electronics and Information Engineering, Tongji University, city=Shanghai, country=China
1 Introduction
@@
In recent years, image anomaly detection techniques have been widely applied in industrial quality detection[1][2][3][4], anomaly segmentation[5], and medical diagnosis scenarios[6][7], aiming to detect abnormal data that are different from normal data within a sample image[8]. Since abnormal prior information is scarce, this task is commonly conducted within the framework of an unsupervised learning paradigm. Consequently, there has been increasing interest among scholars in researching unsupervised anomaly detection, and near-perfect results have been achieved, as evidenced by methods such as Pull&Push[9], PMB-AE[10], and Patchcore[11]. However, due to the limits of those scenarios, most anomaly detection methods currently focus on structural anomaly detection and can only deal with even one object in a single scene at a time, as illustrated in Fig.1(a).
@@
In many usual scenarios such as autonomous driving and surveillance systems, understanding the semantic context of the entire scene and detecting anomalies is essential. Therefore, Bergmann et al.[12] have proposed logical anomalies, which represent more complex abnormalities in the logical relationships between objects within the entire scene. Logical anomalies violate underlying constraints, thereby contravening specific relationships between objects, e.g., a permissible object being present in an invalid location or a required object not being present at all. Specifically, within a scene, anomalies that appear normal at the local level but violate geometric constraints or logical principles when considering global semantics are referred to as global logical anomalies. For example, in Fig.1(b), anomaly-free images of the category breakfast box always contain exactly two tangerines and one nectarine that are always located on the left-hand side of the box. Furthermore, the ratio and relative position of the cereals and the mix of banana chips and almonds on the right-hand side are fixed. An anomaly-free screw bag contains exactly two washers, two nuts, one long screw, and one short screw. Nevertheless, their logical anomalies could involve missing, extra, wrong location, wrong combination, etc.
@@
For the more challenging global logical anomalies, most existing methods tailored for structural anomalies demonstrate catastrophic results, for instance, as illustrated in Fig.1(c) with Patchcore. This is because structural or textural anomalies belong to lower-level anomaly types, which do not necessitate understanding the overall semantics of objects in the scene and only require local knowledge for anomaly detection. However, for higher-level global logical anomalies, relying solely on local perception to ascertain the normalcy of overall semantics is insufficient. Consequently, the performance of existing methods is significantly constrained.
@@
Our previous work[13][14] proposed a reconstruction-based method for detecting multiple object anomalies to address logical anomalies. In [13][14], objects had fixed spatial relationships. We used semi-supervised learning on positive samples to regress object positions in the test samples and reconstruct their normal features. However, exploring new methods is essential to address logical anomalies in more complex scenarios like [12], where objects lack fixed spatial relationships.
@@
To address the above challenges: (1) the presence of multiple key objects in complex scenes, (2) variability in positive sample features, and (3) poor generalization due to dependence on data-driven methods. We propose a novel framework called SAM-LAD, which can be plug-and-play in any scenario without training and even outperforms existing data-driven logical anomaly detection methods. Specifically, we ingeniously leverage the robust object segmentation capability of the Segment Anything Model (SAM)[18] to obtain the positional information of all key objects in the scene by setting segmentation thresholds. Then, we utilize a pre-trained DINOv2[19] backbone network as a feature extractor to extract features and employ nearest neighbor search to find the $k$ most similar normal samples to the query image. Following this, we implement the FeatUp[20] operation to upsample the feature maps and recover their lost spatial information. Combining the upsampled feature maps with the positional information obtained from SAM yields separate feature maps for each object. Subsequently, we propose a Dynamic Channel Graph Attention (DCGA) mechanism to effectively compress each object’s feature map into a single feature vector. Consequently, leveraging the proposed Object Matching Model (OMM), we match the feature vectors of the reference images with those of each feature vector in the query image. Finally, based on the matching results, we propose an Anomaly Measurement Model (AMM), which estimates the feature distribution of individual objects in the query image and the corresponding $k$ matched objects from the $k$ normal samples, thereby calculating the final anomaly score map. Noteworthy, thanks to the introduction of AMM, our framework not only excels in detecting global logical anomalies but also fulfills the requirements for detecting structural anomalies. Our framework is motivated by the following: when humans discern anomalies among multiple objects within intricate scenes, the most straightforward and most efficacious approach involves individually juxtaposing several normal samples with each object in the test sample. Through this meticulous comparison, inference regarding anomalous regions can be inferred without necessitating the establishment of a costly global semantic contextual comprehension within the model. The essence lies in identifying and aligning pivotal objects, thus accomplishing the entirety of the task. We evaluate the proposed framework on multiple commonly used benchmarks, and the experimental results demonstrate that our SAM-LAD achieves state-of-the-art performance. The main contributions of this paper can be summarized as follows:
@@
    We propose the SAM-LAD to address the challenge of detecting logical anomalies. This framework introduces an object-level matching algorithm to determine the correspondence between objects and normal images. Based on the correspondence, a statistical estimator is designed to compute the feature estimation of the object. By analyzing the feature estimation differences between paired objects, we can detect logical and structural anomalies simultaneously, improving the performance of visual anomaly detection models.
@@
    We integrate the visual large model SAM into logical anomaly detection and leveraged its powerful generalization capabilities to achieve zero-shot logical anomaly detection without additional training.
@@
    We conducted experiments on multiple benchmarks, showcasing the state-of-the-art (SoTA) performance of our method.
2 Related Work
2.1 Semantic Segmentation
@@
2.2 Keypoint Matching
Keypoint matching is essential for object recognition, 3D reconstruction, and image stitching in computer vision. The long-standing interest in this problem has led to various approaches. Early methods like SIFT[24] and ORB[25] laid the groundwork with handcrafted keypoint descriptors. These algorithms created robust descriptors invariant to scale, rotation, and partial occlusion, making them practical for many tasks despite sensitivity to image illumination changes. Recently, deep learning techniques revolutionized keypoint matching by learning discriminative features from large datasets. CNNs, like the SuperPoint model[26], have been widely adopted for keypoint detection and matching, learning detection, and descriptor generation end-to-end. Deep learning approaches offer improved robustness and generalization over handcrafted methods. Recent advances use attention mechanisms to improve keypoint matching in challenging environments. Networks like SuperGlue[27] use graph neural networks with self- and cross-attention to establish robust keypoint correspondences. Transformers in keypoint matching improve context understanding, handling complex matching scenarios more proficiently.
@@ 2.2 Keypoint Matching
2.3 Unsupervised Anomaly Detection
@@
Current unsupervised anomaly detection approaches are typically classified into two categories: methods targeting structural anomalies and those targeting logical anomalies. Most current anomaly detection methods target structural anomalies, using robust feature extraction networks to obtain high-level semantic features of a test image. The distance between features of test images and anomaly-free images is calculated, identifying areas with large distances as anomalies. Existing methods typically use deep CNN models like ResNet[28] and EfficientNet[29] for feature extraction. For example, Roth et al.[11] proposed Patchcore to detect anomalies on objects with a concise background. They selected ResNet-50[28] pretrained on ImageNet[30] as the feature extractor. However, these methods struggle with logical anomalies since the local structures within the image appear normal. Therefore, more researchers are focusing on logical anomaly detection. For instance, Zhang et al. proposed DSKD[17], and Batzner introduced EfficientAD[31]. Both methods use the Student-Teacher network, detecting anomalies by comparing teacher and student differences. Using a local-global branching approach, Yao et al. presented GLCF[15], while Bergmann et al. proposed GCAD[12] to understand the logical semantic constraints of the entire image scene. Guo et al. presented THFR[16] based on template-guided reconstruction. These methods have shown promise in detecting certain logical anomalies, such as missing components and unexpected excess. However, their performance drops significantly with more complex logical anomalies, such as misordering, mismatches, and haphazard object arrangements.
@@
The proposed SAM-LAD builds a zero-shot method without strenuous efforts to comprehend the logical semantics of the entire scene. By introducing SAM and proposing OMM and AMM, a global logical anomaly detection system was developed, effectively addressing the limitations of the existing methods mentioned above.
3 Proposed methodology
@@
Figure 2: Pipeline of the proposed framework, which consists of four stages. The first stage is an offline operation, building an anomaly-free template features bank. The second stage is extracting a feature map from a query image. The next stage utilizes SAM to obtain object feature maps further. The last stage involves matching the objects in the query image with those in the reference images one by one. We calculate anomalies within each object and obtain the final anomaly score maps using the obtained matching relationships.
3.1 Architecture Overview
@@
The data flow of the proposed SAM-LAD is depicted in Fig.2, which consists of four stages. 1) On-boarding stage 1 is an offline operation. For all normal images $\mathbb{R}^{3\times H\times W}$, a pre-trained backbone as a feature extractor is first deployed to extract the features from all normal images. Subsequently, the extracted features are compiled and stored within a template features bank. 2) Stage 2 involves extracting features from a query image $I_{q}\in\mathbb{R}^{3\times H\times W}$ using the same feature extractor as stage 1. 3) Stage 3, for the feature maps $f_{q}$ extracted from a query image $I_{q}$, we retrieve its $k$-nearest normal feature maps $(f_{r}^{i})_{i\in{[1,k]}}$ in the template bank and their corresponding images $(I_{r}^{i})_{i\in{[1,k]}}$. We designate the $k$ normal images procured as reference images and in parallel with the query image, input $(I_{q},I_{r}^{i})_{i\in{[1,k]}}$ into the SAM to obtain the individual object mask. Concurrently, the feature maps $(f_{q},f_{r}^{i})_{i\in{[1,k]}}$ are subjected to a FeatUp operation, upsampling to 8$\times$. By combining the individual object masks with the upsampled feature maps, we obtain the object feature maps for both the query and reference images, respectively. 4) In Stage 4, the DCGA module is employed to encode and compress the object feature maps into object descriptor vectors. Subsequently, utilizing the proposed Object Matching Module, the two sets of object descriptor vectors are matched. Based on the matching results, the Anomaly Measurement Module computed the ultimate anomaly score map.
3.2 Feature Extraction and Template Features Bank
@@
The first and second stage of the proposed SAM-LAD is the extraction of feature maps. The same feature maps are later used for FeatUp operation. There are multiple options for extracting features. Recently, Vision Transformer (ViT) has exhibited remarkable performance in anomaly detection-related tasks[32] [33] due to its self-attention mechanism, enabling the model to attend to global information of the entire image when processing image patches. Therefore, we have adopted the pre-trained ViT-type DINOv2-S backbone[19] as the feature extractor for the proposed SAM-LAD. For a given image $I_{x}^{3\times H\times W}$, we denote the extracted feature maps $f_{x}^{C\times H^{{}^{\prime}}\times W^{{}^{\prime}}}$:
@@
At initialization, we execute offline operations to construct a template features bank =$\mathcal{B}=\left\{B_{1},\cdots,B_{i},\cdots,B_{Z},\right\}$ using $F$ to extract $Z$ normal images in the all normal set $\mathbb{R}^{3\times H\times W}$, where $B_{i}$ represents the template feature map of the $i$-th normal sample. At inference, only the feature maps of the query image are extracted, and we could reduce the template features bank using coreset subsampling method[34][35] to reduce inference time and memory usage.
3.3 Image-level Nearest Neighbor Search